AP 10-3 Using the Effective-Interest Method with Explanation of bond Premium and Discussion of Management strategy.

On March 1,2018, Carter Corporation issued $15,000,000 in bonds that mature in 10 years. The bonds have a coupon rate of 6.3 percent and pay interest on March 1 and September 1. When the bonds were sold, the market rate of interest was 6 percent. Carter uses the effective-interest method to amortize bond discount or premium. By December 31 , 2018, the market interest rate had increased to 7 percent.

REQUIRED:

1. Record the issuance of the bond on March 1, 2018.

2. Compute the present value of the difference between the interest paid each six months ($472,500) and the interest demanded by the market ($ 15 million x 6% x 6/12 = $ 450,000). Use the market rate of interest and the 10-year life of the bond in your present value computation. What does this amount represent? Explain.

3. Record the payment of interest on September 1, 2018.

4. Record the adjusting entry for accrued interest on December 31, 2018.

5. Why does the interest expense change each year when the effective-interest method is used?

6. Compute the present value of Carter's bonds, assuming that they had a 7-year life instead of a 10-year life. Compare this amount to the carrying amount of the bond at March 1, 2021. What does this comparison demonstrate?

7. Determine the impact of the transactions at year-end on the debt-to-equity ratio and the times interest earned ratio.
Answer #1

1. Cash Ac Dr $15,000,000

To Bonds Payable $15,000,000

(Being bonds issued at face value)

2.

Difference between interest paid:(472500-450000)=$22,500

PV Value of Difference:22500-21844.66=$655.34

*22500/1.03=21844.66

3. Bond Interest Expenses ac Dr.472,500

To cash ac 472500

4. Bond Interest Expenses Ac Dr. 472,500

To Interest Payable Ac 472,500